Options trade on the common stock of Taz, Inc. that have a strike price of $50.00 and a premium of $2.50. In each of the next four parts, calculate the net profit (or loss) on the option position, where net profit includes the premium in the calculation. Note: Negative responses should be placed with a preceding negative sign (e.g. -4.50) and not with parentheses (e.g. (4.50)). Part 1: Calculate the net profit or loss from BUYING a CALL option on Taz if at the time of expiration the price per share of Taz is $52.75. $ Part 2: Calculate the net profit or loss from WRITING a CALL option on Taz if at the time of expiration the price per share of Taz is $52.75. $ Part 3: Calcuate the net profit or loss from BUYING a PUT option on Taz if at the time of expiration the price per share of Taz is $52.75. $ Part 4: Calcuate the net profit or loss from WRITING a PUT option on Taz if at the time of expiration the price per share of Taz is $52.75. $
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General aCopy B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 conditional Format . Cell Insert Delete Format Paste Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. Styles ▼ ㆆ ▼ Clipboard Font Alignment Number Cells IJ69 Formula Bar IH IK IM IN IO Ia IR Is IT IU 69 70 71 72 73 74 75 76 TAZ INC STRIKE PRICE PREMIUM 2.5 PART 1 BUYING A CALL PRICE ON EXPIRY-52.75 AS PRICE ON EXPIRY>STRIKE PRICE, CALL SHOULD BE EXERCISED NET PROFIT PAYOFF PREMIUM PAID NET PROFIT (PRICE ON EXPIRY-STRIKE PRICE) - PREMIUM PAID NET PROFIT= (52.75-50)-2.50 = 0.25 78 79 80 81 82 83 84 85 86 87 4 ‘ ト PART 2 WRITING A CALL: WE HAVE SOLD CALL OPTIONS, WE ARE SELLER, WE WILL RECEIVE PREMIUM PRICE ON EXPIRY = 52.75 AS PRICE ON EXPIRY> STRIKE PRICE, BUYER OF CALL OPTION WILL EXERCISE THAT MEANS WE HAVE NEGATIVE PAY OFF NET PROFIT PAYOFFPREMIUM RECEIVED NET PROFIT (50-52.75) + 2.50-0.25 OCF CASH FLOW STAT N -FCF -ARM MSI confidence!NDEX CURRENCY LOAN CROSS OVER HEDGING MONEY MARKET GM HPR AM TAX GROWTH SUST PREF AND DIV POLICY IHome nert Page Layout Formulas Data Review View dd-Ins Σ AutoSum Cut ー E ゴWrap Text Sort &Find & Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ ta copy. B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 2 ClearFe Select Edting Paste Format Painter Cells Number Alignment Clipboard IO IP Ia IR IS IT IU IH IJ IK IM IN 87 BUYING A PUT PRICE ON EXPIRY 52.75 AS PRICE ON EXPIRY> STRIKE PRICE, PUT SHOULD NOT BE EXERCISED, PAYOFF WILL BE ZERO NET PROFIT PAYOFF-PREMIUM PAID NET PROFIT = (0)-2.50 = PART 3 89 90 91 92 93 94 95 96 97 98 -2.50 PART 4WRITING A PUT: WE HAVE SOLD PUT OPTIONS, WE ARE SELLER, WE WILL RECEIVE PREMIUM PRICE ON EXPIRY = 52.75 AS PRICE ON EXPIRY>STRIKE PRICE, BUYER OF PUT OPTION WILL NOT EXERCISE THAT MEANS WE HAVE ZERO PAYOFF NET PROFIT PAYOFF + PREMIUM RECEIVED NET PROFIT = 0 + 2.50 = 2.50 100 101 102 103 104 105 HEDGING MONEY MARKET GM HPR AM TAXGROWTH SUSTPREF AND DIV POLICY , OCF CASH FLOW STATN ARM, MSI confidenceINDEX CURRENCYLOANCROSS OVER 09-01-2019
Answer #2

THIS COPIED FROM MY SOLUTION. THIS IS MY EXCEL FILE

answered by: pravinmandora